Find the z-scores that bound the middle 15% of the standard normal distribution (give anser to 2 decimal places)

Smaller value -

Larger value -
Answer #1

Solution :

Given that,

Using standard normal table,

P(-z < Z < z) = 15%
P(Z < z) - P(Z < z) = 0. 15
2P(Z < z) - 1 = 0. 15
2P(Z < z ) = 1 + 0.15
2P(Z < z) = 1.15
P(Z < z) = 1.15 / 2
P(Z < z) = 0.575
z =0.189 and z = - 0.189

Smaller value = - 0.19

Larger value =0.19